OpenStudy (anonymous):

An arithmetic sequence is represented by the explicit formula A(n) = 2 + 9(n - 1). What is the recursive formula?

OpenStudy (anonymous):

A. A(n) = A(n - 1) + 2 B. A(n) = A(n - 1) + 9 C. A(n) = A(n - 1) - 9 D. A(n) = A(n - 1) - 2